# Multimodal Latent Diffusion Model for Complex Sewing Pattern Generation

### v1

@@ 1 Introduction
Clothes play a pivotal role in shaping human aesthetics and physique, where appropriate clothes can beautify their overall appearance and highlight human physical attributes. Therefore, garment design has always been a crucial component that significantly impacts both digital character creation and real-life human visual presentation. In recent years, many garment generation methods [52, 29, 26, 20, 33, 27, 53, 4, 42, 43, 37, 9, 61, 71, 25] have emerged to generate desired garments for users through various conditions.
@@
Typically, garment generation can be classified into 2D and 3D methods. The 2D generation methods [4, 42, 43, 25] can produce visually appealing results but cannot maintain consistency across different views, failing to drape on human bodies. Therefore, many recent works are focusing on 3D cloth generation [52, 29]. Although these methods can generate high-quality mesh or neural field, they pose the challenge of clipping between clothes and bodies when draping clothes onto the human body, and they are incompatible with the digital garment production pipeline. Meanwhile, the sewing pattern is a more widely used representation for garments in the industry because it facilitates both physical simulation and animation in CG-friendly fashions [3, 7]. Current methods for sewing pattern generation [26, 20, 33] achieve fantastic garments generation but fall short in designing complex features, such as gathers and darts. Besides, these methods typically ignore human body shapes, preventing the creation of the made-to-measure garment. Apart from learning-based methods, parametric garment design tool [27] allows users to model complex garments and considers their relations with human body shapes. Nonetheless, this tool requires pre-defined templates and a delicate selection of control parameters, which require professional knowledge of garment designs, limiting its widespread promotion. In summary, there are two main challenges in suitable sewing pattern generation: 1) Designing a more general representation for complex designs of sewing patterns, and 2) Enabling control over garment details and ensuring the garments are body-suited.
@@
To address these issues, we design a novel architecture named SewingLDM for generating complex sewing patterns under the control of texts, body shapes, and garment sketches. To represent complex designs of sewing patterns, we especially design an extended representation to encompass intricate types of edges and attachments of garments, enabling more general and complex garment learning. Subsequently, we train an auto-encoder model to compress the representation into a compact latent space for easier training while maintaining high reconstruction quality. To achieve multi-modal controlled and body-aware sewing pattern generation, we design a two-step training strategy to introduce the control signals into the latent diffusion model. In the first step, we train the diffusion model under the condition of texts, serving as a coarse fundamental model for additional control signal injection. In the second step, we further embed the knowledge of sketches and body shapes into the diffusion model by fusing the features after the first block and fine-tuning the output layers within the attention module, providing additional control of garment details and ensuring the generated garments fit various body shapes. Based on the proposed framework, SewingLDM can generate complex garments that fit various body shapes and align with user-provided text descriptions or garment sketches.
@@
  • We design an extended representation to cover complex sewing patterns and compress it into a compact latent space for easier training of the generation model.
@@ 2 Related Work
Multimodal-guided 3D Generation. The advancements in large models [49, 51, 1] have encouraged the emergence of recent 3D generation models [57, 23, 21, 41, 11, 34, 68], entering a new era of 3D generation. Some models [47, 55, 65, 8, 40, 37] focus on generating implicit neural radiance fields corresponding to the text description, whereas others [56, 11] extend their ability to generating 3D meshes with BRDF materials. Especially, some models [52, 53, 29, 37] dive into human clothes generation. Garment3DGen [52] can generate textured 3D mesh from images or text conditions by adjusting template meshes. GarmentDreamer [29] utilizes 3D Gaussian Splatting (GS) [24] as guidance to create 3D garment meshes from textual prompts. WordRobe [53] leverages unsigned distance field (UDF) [18] to represent 3D garments and generate 3D garment meshes under text guidance. GarVerseLOD [37] proposes a hierarchical framework to recover different levels of garment details through single images. Although these methods can generate visual-appealing garment meshes, their compatibility with CG pipelines remains a challenge, hindering seamless integration into modern industry workflows. In contrast, our method aims at sewing pattern generation, facilitating utilization in CG processes and daily clothing factories.
@@
3D Sewing Pattern Modeling. Existing fashion CAD software tools, such as Clo3D [14] and Marvelous Designer [15], allow users to edit sewing patterns and simulate desired cloth outcomes. While these methods integrate the most advanced garment design, they heavily rely on artists to manually draw and adjust the shapes of sewing patterns, requiring a substantial of professional manual processing. Consequently, ongoing studies are focusing on automating the adjustment of sewing patterns [6, 58, 38, 32, 60, 46, 16], reconstructing sewing patterns [22, 62, 19, 12, 5, 26, 33], and assisting with complex garment design [30, 59, 17, 13, 31]. Recent studies [33, 20, 26, 27] on sewing patterns start to focus on autonomously generating diverse sewing patterns through different conditions rather than merely adjusting or producing a single garment. One of the recent SOTA methods, DressCode [20], first generates garments through natural language and yields visual-appealing appearances. However, the capability of DressCode is limited in modeling complex sewing patterns, and furthermore, it does not consider the relation between garments and body shapes, difficult to drape on various bodies. Another typical work, parametric sewing pattern [27], can control complex sewing patterns, while it requires predefined templates and a delicate selection of different control scale values, which is not user-friendly. Different from these methods, our SewingLDM not only has the ability to represent complex sewing patterns, but can also generate sewing patterns based on multi-modal intuitive conditions, i.e., natural language, garment sketches, and body shape. These capabilities enable easily creating tailored garments that conform precisely to individual body shapes.
@@ 3 Method
To generate garments suited for various humans, we introduce SewingLDM, a latent-based diffusion model, to create complex 3D sewing patterns, conditioned by personalized body shapes, text prompts, and garment sketches. We first review the original sewing pattern representation [26] (Sec. 3.1) and then we improve it with special designs to cover complex sewing patterns (Sec. 3.2), as illustrated in Fig. 2. Subsequently, we compress the sewing pattern representation into a compact latent space for easier training of the generation model and reducing computational resources in Fig. 3 (Sec. 3.3). Finally, we train a latent diffusion model under multi-modal conditions through a two-stage training strategy, as illustrated in Fig. 4 (Sec. 3.4). Based on the proposed framework, SewingLDM can generate complex garments based on the body shape and align with the user-provided text description or garment sketches.
@@ 3.2 Extended Representation for Sewing Pattern
Figure 4: Multimodal latent diffusion model. After training the text-guided diffusion model, we fuse the features of sketches and body shapes and normalize them into the diffusion model, with fine-tuning minimal parameters of the diffusion model. The trained network parameters are depicted in orange, while the frozen parameters are shown in purple. The output latent is then quantized into a designed latent space and serves as the input of the decoder to yield all edge lines. Edge lines connect from beginning to end to form panels, placed on the corresponding body regions. Finally, we can get suited garments through the modern CG pipeline.
@@
Preprocessing. Before input to the neural networks, the vector representation needs to be the same size for all data during training. For each edge $E_{i,j}$, we concatenate all extended parameters and append with rotation $R_{i}$ and translation $T_{i}$ of panel $P_{i}$ to form the high dimensional edge feature $E^{f}_{i,j}$. Furthermore, we design a binary flag $\{E^{m}_{i,j}\in\{0,1\}\}_{j=1}^{N_{i}}$ to denote the existence of each edge. All features are concatenated to form a 29-dimensional vector for each edge feature $E^{f}_{i,j}$, represented as follows:
@@
Then, all edge features $\{E^{f}_{i,j}\}_{j=1}^{N_{i}}$ are concatenated and padded with 0 to max edge number to get the representation of panels. Further, we concatenate all panels and pad with 0 to max panel number to get the representation of sewing pattern $F$, in the shape of $(max(N_{p})\times max(N_{i}),29)$. Before input to the neural network, all continuous values are standardized, and all binary flags are transformed into $\{-1,1\}$.
@@ 3.3 Compact Latent for Sewing Pattern
The vector representation $F$ inevitably incorporates redundant information as the panel and edge numbers increase, preventing generation models from learning the distribution of $F$. As indicated by the recent compression methods [70, 39, 64], it is necessary to compress $F$ into a compact latent space and maintain the reconstruction quality. Following this idea, we train an auto-encoder to compress and quantize the $F$ into a latent space where each dimension is bounded within the range $\left[-1,1\right]$, as illustrated in Fig. 3. The sewing pattern representation $F$ is encoded to $z$ by the encoder $\mathcal{E}$ and quantized to $\hat{z}$ in the constrained latent space, subsequently reconstructed by the decoder $\mathcal{D}$. The process can be represented as:
@@
where $n$ is an integer used to modify the spacing between each $\hat{z}$ in the latent space.
For training the encoder $\mathcal{E}$ and decoder $\mathcal{D}$, we combine the loss in previous works [26, 39] with additional binary cross-entropy loss $\mathcal{L}_{\mathrm{BCE}}$ to constrain the newly incorporated binary flags:
@@ 3.4 Multimodal Conditions of Diffusion Model
The text-guided latent diffusion model serves as a fundamental model for extensive multi-modal conditions injection. For the injection of body shapes and garment sketches, a naive idea is to inject them through two ControlNet [67] branches. While sketches will change along with body shape, e.g., sketches will get wider when bodies grow fatter. We propose to first use embedders to extract the feature from sketches and body shapes, and then simply concatenate them together, and input it into a light transformer layer. During the light transformer layer, the features of sketches and body shapes can be thoroughly fused and get the relation between each other through self-attention modules and output as $F_{bs}$. Then we normalize mean $\bm{\mu}_{bs}$ and variance $\bm{\sigma}_{bs}$ of $\bm{F}_{bs}$ into the same mean $\bm{\mu}_{z}$ and variance $\bm{\sigma}_{z}$ with the latent features $\bm{F}_{z}$ and add them together during the middle block of the diffusion model.
@@
where $\epsilon$ is a small constant for numerical stability. After the normalization, we assume the new $\hat{\bm{F}}_{z}$ is similar to $\bm{F}_{z}$, which does not need to retrain the whole diffusion model in the second stage. We only fine-tune the output layer of the attention modules in each DiT block to transform the normalized features into the desired distribution. After two-stage training, our generation model can precisely follow the text guidance and sketches under various human shapes, enabling more body-suited and detailed controlled garment generation for individuals.
@@ 4.1 Experiment Setup
Dataset. To train a generation model under the condition of texts or sketches, it is essential to acquire the corresponding paired data. We extend the current dataset [28] with additional textual annotations and garment sketches. The dataset [28] consists of 120,000 sewing patterns, covering a variety of clothing styles for different body types. Sewing patterns in [28] consider the relationship between various body shapes and garments, resulting in garments that are well-tailored to individual body types. Building on [28], we annotate each garment with text prompts according to its design parameters file, resulting in detailed text annotations. However, relying solely on textual descriptions may not precisely dictate garment shapes, potentially yielding undesirable outputs. To enhance control over the generation, we propose to generate more rich annotations like sketches. For each garment, we utilize PiDiNet [54], a pre-trained edge detection network, to extract garment sketches, thereby enriching the design details of the garment.
@@ 4.2 Qualitative Comparison
Comparison of Sewing Pattern Generation Methods. We also compare our method with current SOTA sewing pattern generation methods, i.e., DressCode [20] and Sewformer [33]. DressCode and our SewingLDM are fed with the same text descriptions, and we use ControlNet to generate the corresponding images as the input of Sewformer [33]. Additionally, the sketches used to generate images are utilized in our SewingLDM. Furthermore, the generative sewing patterns are draped onto two different body shapes to validate the effectiveness of body-aware garment generation. As illustrated in Fig. 6, both Sewformer and DressCode fail to generate textual-aligned garments due to the complex garment descriptions. Moreover, they can not be worn to diverse body shapes, sliding from the body or just failing to simulate the results. In contrast, our method can generate complex sewing patterns, e.g., mermaid skirt hem, one-shoulder gown, and circle neckline, and fit to various body shapes, which provides a more user-friendly approach to getting the desired made-to-measure garments. The results demonstrate the superiority of our method in complex garment design and various body adaptability.
4.3 Quantitative Comparison
@@
Table 1: Quantitative Comparison. We compare the generation efficiency and the average clothes-to-body distance. Further, we conduct a comprehensive user study to judge the superiority of different methods. All metrics show our method generation superior results than other methods.
Besides qualitative comparisons, we also perform quantitative comparisons with these SOTA methods, evaluating aspects including generation efficiency, clothes-to-body distance, and user study. The clothes-to-body distance is to indicate whether the clothes are close-fitted to body shapes calculated by averaging the minimum distance from each point in garments to human bodies. Besides, we also perform a user study to further assess the quality of garment generation. We take 10 text prompts to generate diverse garments and render the generated results draping on different body shapes for each method. Then we ask 30 users to give a value of these rendering results with comprehensive consideration for two aspects: 1) consistency with text descriptions; and 2) well-fitting with human bodies. As illustrated in Tab. 1, the preference results demonstrate a notable superiority of our method over SOTA approaches in both aspects, highlighting in generating garments that are both well-suited to various bodies and exhibit high fidelity aligned with text descriptions.

### v2

@@ 1 Introduction
Clothes play a pivotal role in shaping human aesthetics and physique, where appropriate clothes can beautify their overall appearance and highlight human physical attributes. Therefore, garment design has always been a crucial component that significantly impacts both digital character creation and real-life human visual presentation. In recent years, many garment generation methods [55, 30, 27, 21, 34, 28, 56, 4, 44, 45, 39, 10, 65, 75, 26] have emerged to generate desired garments for users under various conditions.
@@
Typically, garment generation can be classified into 2D and 3D methods. The 2D generation methods [4, 44, 45, 26] can produce visually appealing results but cannot maintain consistency across different views, failing to drape on human bodies. Therefore, many recent works are focusing on 3D cloth generation [55, 30]. Although these methods can generate high-quality meshes or neural fields, they pose the challenge of clipping between clothes and bodies when draping clothes onto the human body, and they are incompatible with the digital garment production pipeline. Meanwhile, the sewing pattern is a more widely used representation for garments in the industry because it facilitates both physical simulation and animation in CG-friendly fashions [3, 8]. Previous methods for sewing pattern generation [27, 21, 34] achieve fantastic garment generation but fall short in designing complex features. Most recently, concurrent works [7, 46, 76] introduce complex features with reduced tokenization, leading to significant advancements in garment generation. However, these methods typically ignore human body shapes, preventing the creation of made-to-measure garments. Apart from learning-based methods, the parametric garment design tool [28] allows users to model complex garments and considers relations with body shapes. Nonetheless, this tool requires pre-defined templates and detailed body measurements, and then with a delicate selection of control parameters to generate the garments, which require professional knowledge of garment designs. In summary, there are two main challenges in suitable sewing pattern generation: 1) Designing a general representation for complex sewing patterns, and 2) Enabling control over garment details and ensuring garments are body-suited.
@@
To address these issues, we design a novel architecture named SewingLDM for generating complex sewing patterns under the control of texts, body shapes, and garment sketches. To represent complex designs of sewing patterns, we especially design an extended representation to encompass intricate types of edges and attachments of garments, enabling more general and complex garment learning. Subsequently, we train an auto-encoder model to compress the representation into a compact latent space, which not only enables efficient training of complex sewing pattern generation with fixed memory consumption but also ensures scalability for future applications involving more intricate garment designs. To achieve multi-modal controlled and body-aware sewing pattern generation, we design a two-step training strategy to introduce the control signals into the latent diffusion model. In the first step, we train the diffusion model under the condition of texts, serving as a coarse fundamental model for additional control signal injection. In the second step, we further embed the knowledge of sketches and body shapes into the diffusion model by fusing the features after the first block and fine-tuning the output layers within the attention module, providing additional control of garment details and ensuring the generated garments fit various body shapes end to end without the need of delicate body measurements. Based on the proposed framework, SewingLDM can generate complex garments that fit various body shapes and align with user-provided text descriptions or garment sketches.
@@
  • We design an extended representation to cover complex sewing patterns and compress it into a compact latent space enabling the training of complex sewing patterns generation and maintaining high reconstruction quality.
@@ 2 Related Work
Multimodal-guided 3D Generation. The advancements in large models [52, 54, 1] have encouraged the emergence of recent 3D generation models [61, 24, 22, 43, 12, 35, 72, 58], entering a new era of 3D generation. Some models [50, 59, 69, 9, 42, 39] focus on generating implicit neural radiance fields corresponding to the text description, whereas others [60, 12, 36] extend their ability to generate 3D meshes with BRDF materials. Especially, some models [55, 56, 30, 39] dive into human clothes generation. Garment3DGen [55] can generate textured 3D mesh from images or text conditions by adjusting template meshes. GarmentDreamer [30] utilizes 3D Gaussian Splatting (GS) [25] as guidance to create 3D garment meshes from textual prompts. WordRobe [56] leverages unsigned distance field (UDF) [19] to represent 3D garments and generate 3D garment meshes under text guidance. GarVerseLOD [39] proposes a hierarchical framework to recover different levels of garment details through single images. Although these methods can generate visual-appealing garment meshes, their compatibility with CG pipelines remains a challenge, hindering seamless integration into modern industry workflows. In contrast, our method aims at sewing pattern generation, facilitating utilization in CG processes.
@@
3D Sewing Pattern Modeling. Existing fashion CAD software tools, such as Clo3D [15] and Marvelous Designer [16], allow users to edit sewing patterns and simulate desired cloth outcomes. While these methods integrate the most advanced garment design, they heavily rely on artists to manually draw and adjust the shapes of sewing patterns, requiring a substantial of professional manual processing. Consequently, ongoing studies are focusing on automating the adjustment of sewing patterns [6, 62, 40, 33, 64, 49, 17], reconstructing sewing patterns [23, 66, 20, 13, 5, 27, 34], and assisting with complex garment design [31, 63, 18, 14, 32]. Recent studies [34, 21, 27, 28, 7, 46, 76] on sewing patterns start to focus on autonomously generating diverse sewing patterns through different conditions rather than merely adjusting or producing a single garment. One of the recent SOTA methods, DressCode [21], first generates garments through natural language and yields visual-appealing appearances. However, the capability of DressCode is limited in modeling complex sewing patterns, and furthermore, it does not consider the relation between garments and body shapes, difficult to drape on various bodies. Another typical work, parametric sewing pattern [28], can control complex sewing patterns, while it requires predefined templates and a delicate selection of different control scale values, which is not user-friendly. Different from these methods, our SewingLDM not only has the ability to represent complex sewing patterns, but can also generate sewing patterns based on multi-modal intuitive conditions, i.e., natural language, garment sketches, and body shape. These capabilities enable easily creating tailored garments that conform precisely to individual body shapes.
@@ 3 Method
To generate garments suited for various humans, we introduce SewingLDM, a latent-based diffusion model, to create complex 3D sewing patterns, conditioned by personalized body shapes, text prompts, and garment sketches. We first review the original sewing pattern representation [27] (Sec. 3.1) and then we improve it with special designs to cover complex sewing patterns (Sec. 3.2), as illustrated in Fig. 2. Subsequently, we compress the sewing pattern representation into a compact latent space enabling the generation of complex sewing patterns with centimeter-level precision in Fig. 3 (Sec. 3.3). Finally, we train a latent diffusion model under multi-modal conditions through a two-stage training strategy, as illustrated in Fig. 4 (Sec. 3.4). Based on the proposed framework, SewingLDM can generate complex garments based on the body shape and align with the user-provided text description or garment sketches.
Figure 4: Multimodal latent diffusion model. After training the text-guided diffusion model, we fuse the features of sketches and body shapes and normalize them into the diffusion model, with minimal fine-tuning parameters of the diffusion model. The trained network parameters are depicted in orange, while the frozen parameters are shown in purple. The output latent is then quantized into a designed latent space and serves as the input of the decoder to yield all edge lines. Edge lines connect from beginning to end to form panels, placed on the corresponding body regions. Finally, we can get suited garments through the modern CG pipeline.
@@ 3.2 Extended Representation for Sewing Pattern
Preprocessing. Before input to the neural networks, the vector representation needs to be the same size for all data during training. For each edge $E_{i,j}$, we concatenate all extended parameters and append with rotation $R_{i}$ and translation $T_{i}$ of panel $P_{i}$ to form the high-dimensional edge feature $E^{f}_{i,j}$. Furthermore, we design a binary flag $\{E^{m}_{i,j}\in\{0,1\}\}_{j=1}^{N_{i}}$ to denote the existence of each edge. All features are concatenated to form a 29-dimensional vector for each edge feature $E^{f}_{i,j}$, represented as follows:
@@
Then, all edge features $\{E^{f}_{i,j}\}_{j=1}^{N_{i}}$ are concatenated and padded with 0 to max edge number and max panel number to get the representation of sewing pattern $F$, in the shape of $(max(N_{p})\times max(N_{i}),29)$. Before input to the neural network, all continuous values are standardized, and all binary flags are transformed into $\{-1,1\}$.
@@ 3.3 Compact Latent for Sewing Pattern
The vector representation $F$ inevitably incorporates redundant information as the panel and edge numbers increase, preventing generation models from learning the distribution of $F$. Using the tokenization in the previous method [21] will lead to excessive GPU consumption, exceeding hardware limitations. As indicated by the recent compression methods [74, 41, 68], it is necessary to compress $F$ into a compact latent space and maintain the reconstruction quality. Following this idea, we train an auto-encoder to compress and quantize the $F$ into a latent space where each dimension is bounded within the range $\left[-1,1\right]$, as illustrated in Fig. 3. The sewing pattern representation $F$ is encoded to $z$ by the encoder $\mathcal{E}$ and quantized to $\hat{z}$ in the constrained latent space, subsequently reconstructed by the decoder $\mathcal{D}$. The process can be represented as:
@@
where $n$ is an integer used to modify the spacing between each $\hat{z}$ in the latent space. For training the encoder $\mathcal{E}$ and decoder $\mathcal{D}$, we combine the loss in previous works [27, 41] with additional binary cross-entropy loss $\mathcal{L}_{\mathrm{BCE}}$ to constrain the newly incorporated binary flags:
@@ 3.4 Multimodal Conditions of Diffusion Model
The text-guided latent diffusion model serves as a fundamental model for extensive multi-modal conditions injection. For the injection of body shapes and garment sketches, a naive idea is to inject them through two ControlNet [71] branches. While sketches will change along with body shape, e.g., sketches will get wider when bodies grow fatter. We propose to first use embedders to extract the feature from sketches and body shapes, and then simply concatenate them together, and input them into a light transformer layer. During the light transformer layer, the features of sketches and body shapes can be thoroughly fused and get the relation between each other through self-attention modules and output as $\bm{F}_{bs}$. Then we normalize mean $\bm{\mu}_{bs}$ and variance $\bm{\sigma}_{bs}$ of $\bm{F}_{bs}$ into the same mean $\bm{\mu}_{z}$ and variance $\bm{\sigma}_{z}$ with the latent features $\bm{F}_{z}$, serving as a residual to control the output. The new $\hat{\bm{F}}_{z}$ is represented as below:
@@
where $\epsilon$ is a small constant for numerical stability. After the normalization, we assume the newly added residual is similar to $\bm{F}_{z}$, which does not need to retrain the whole diffusion model in the second stage. We only fine-tune the output layer of the attention modules in each DiT block to transform the normalized features into the desired distribution. After two-stage training, our generation model can precisely follow the text guidance and sketches under various human shapes, enabling more body-suited and detailed controlled garment generation for individuals.
@@ 4.1 Experiment Setup
Dataset. To train a generation model under the condition of texts or sketches, it is essential to acquire the corresponding paired data. We extend the current dataset [29] with additional textual annotations and garment sketches. The dataset [29] consists of 120,000 sewing patterns, covering a variety of clothing styles for different body types. Sewing patterns in [29] consider the relationship between various body shapes and garments, resulting in garments that are well-tailored to individual body types. Building on [29], we annotate each garment with text prompts according to its design parameters file and refine it with GPT-4 [1], resulting in detailed text annotations. However, relying solely on textual descriptions may not precisely dictate garment shapes, potentially yielding undesirable outputs. To enhance control over the generation, we propose to generate richer annotations like sketches. For each garment, we utilize PiDiNet [57], a pre-trained edge detection network, to extract garment sketches, thereby enriching the design details of the garment. More annotation details are represented in the supplementary material.
@@ 4.2 Qualitative Comparison
Comparison of Sewing Pattern Generation Methods. We also compare our method with current SOTA sewing pattern generation methods, i.e., DressCode [21] and Sewformer [34]. To ensure a fair comparison, we directly utilize the in-the-wild images chosen from the Multimodal Garment Designer [4] and transfer into the corresponding conditions of DressCode and Sewformer. Moreover, since Sewformer and DressCode are only trained with the average SMPL body shapes, we drape the generative sewing patterns onto the average body shape and another body shape to validate the effectiveness of body-aware garment generation. As illustrated in Fig. 6, both Sewformer and DressCode fail to generate textual-aligned garments due to the complex garment descriptions. Moreover, since these baseline methods do not account for various bodies and generate attachments to any edge of sewing patterns, they can not be worn on diverse body shapes, only on the average body shape, sliding from the body or just failing to simulate the results. In contrast, our method uses a compact latent space to represent complex sewing patterns with centimeter-level precision, which greatly improves the generation ability on complex sewing pattern generation, e.g., fitted shirts, one-shoulder gowns, and square necklines. Moreover, due to the body shape conditions, our method can fit various body shapes, which provides a more user-friendly approach to getting the desired made-to-measure garments. The results demonstrate the superiority of our method in complex garment design and various body adaptability. We also provide concurrent work comparison and more in-the-wild and in-domain conditions in the supplementary material.
@@
Table 1: Quantitative Comparison. We compare the generation efficiency, the average clothes-to-body distance, and users’ evaluation. All metrics show that our method generated superior results.
Method Panel L2()($\downarrow$)( ↓ ) Panel Acc()($\uparrow$)( ↑ ) Edge Acc()($\uparrow$)( ↑ ) Rot L2()($\downarrow$)( ↓ ) Transl L2()($\downarrow$)( ↓ ) Stitch Acc()($\uparrow$)( ↑ ) Failure Rate()($\downarrow$)( ↓ )
SewFormer* 12.3 79.4 44.7 .0400 4.5 2.8 4.3%
AE (Ours) 0.64 99.8 88.5 .0004 0.6 90.8 0
SewingLDM 3.13 97.8 82.7 .0043 1.2 84.2 0
Table 2: Reconstruction metrics. We report the reconstruction performance on the test set of the GarmentCode dataset [29]. Our auto-encoder (AE) can reconstruct the sewing pattern with centimeter-level precision compliant with industrial standards.
4.3 Quantitative Comparison
Besides qualitative comparisons, we also perform quantitative comparisons with these SOTA methods, evaluating aspects including generation efficiency, clothes-to-body distance, and user study. The clothes-to-body distance is an essential metric for body-aware garment generation, indicating whether the clothes are close-fitted to body shapes calculated by averaging the minimum distance from each garment point to human bodies. Besides, we also perform a user study to further assess the quality of garment generation. We take 10 text prompts to generate diverse garments and render the generated results, draping on different body shapes for each method. Then we ask 30 users to give a value of these rendering results with comprehensive consideration for two aspects: 1) consistency with text descriptions; and 2) well-fitting with human bodies. As illustrated in Tab. 1, the preference results demonstrate a notable superiority of our method over SOTA approaches in both aspects, highlighting in generating garments that are both well-suited to various bodies and exhibit high fidelity aligned with text descriptions. Furthermore, we also evaluate sewing pattern reconstruction metrics as proposed in SewFormer [34], including: 1) Panel L2, Rot L2, and Transl L2 (average edge, rotation, and translation L2 distances between predicted and ground-truth patterns); 2) Panel Accuracy, Edge Accuracy, Stitch Accuracy (accuracy of panel count, edge prediction per panel, stitch prediction); and 3) Failure Rate (simulation failure percentage). All L2 metrics are in centimeters, except rotation in radians. To ensure a fair evaluation, SewFormer* is fine-tuned on the GarmentCode dataset until its validation loss no longer improves. Dresscode, which requires over 30k tokens for complex patterns and exceeds current GPU limitations, unabling to train on the GarmentCode dataset, was evaluated only for failure rate (11.4%) using its open-source code and checkpoint. As shown in Tab. 2, our method achieved superior performance compared with baselines on complex sewing pattern generation as our latent space is well-designed and suitable for centimeter-level garments compression and our model takes the body shapes into account.